OpenStudy (anonymous):

s(x,y) = x^3 lny + 4y^2e^x find the indicated function Syy (x, y)

OpenStudy (anonymous):

I think it's [-X^3/y^(2)] +[8e^(x)]

OpenStudy (anonymous):

Yes. it is, but how ? could you show me?

OpenStudy (anonymous):

yeah but give me like 10 15 minutes i'm going to grab something to eat

OpenStudy (anonymous):

ok, well you understand that when taking a Sy x is held constant.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay cool then what's the derivative of lny?

OpenStudy (anonymous):

1/y

OpenStudy (anonymous):

?

OpenStudy (anonymous):

x^3 is constant and that would stay on the top correct? next, What is the derivative of 4y^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8y

OpenStudy (anonymous):

e^x is constant so, what is Sy?

OpenStudy (anonymous):

8e^x

OpenStudy (anonymous):

not exactly, the answers that we found previously will be Sy. What I mean by that it Sy :x^3/y +8ye^x. These are the answers you gave me. Do you see how it equals that? If not, we can approach it differently.

OpenStudy (anonymous):

So Ssubyy = -x^3/y^2 + 8e^x?

OpenStudy (anonymous):

remember "x" is constant so if you want plug in 7 for x just so you get use to the idea

OpenStudy (anonymous):

sorry yea that's the answer

OpenStudy (anonymous):

Yes.. why are you plugging in 7? I understand Sy = x^3/y + 8ye^x.. I think my question was to find Syy (x,y) =?

OpenStudy (anonymous):

oh then you just do a second partial it's the same idea as doing a second derivative and the "7" thing is just case if haven't grasped the idea that the other variable is held constant.

OpenStudy (anonymous):

Ic

OpenStudy (anonymous):

Are you in Calc 3 or DE?

OpenStudy (anonymous):

No.. just very basic business calc

OpenStudy (anonymous):

cant imagine doing the regular calc :(

OpenStudy (anonymous):

you have to do partials?

OpenStudy (anonymous):

Yes.. now to find the local extreme for max and min in 3D

OpenStudy (anonymous):

I have another one.. I just can't do figure out should I post it separate?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

I just fan you.. thanks for your help

OpenStudy (anonymous):

I tried to post it.. but it got stuck f(x,y) = x/y = y/x find fxx, fxy, and fyy